### 29 Jan 2024 v1

# Attention-based Reinforcement Learning for Combinatorial Optimization: Application to Job Shop Scheduling Problem

Abstract
Job shop scheduling problems are one of the most important and challenging combinatorial optimization problems that have been tackled mainly by exact or approximate solution approaches. However, finding an exact solution can be infeasible for real-world problems, and even with an approximate solution approach, it can require a prohibitive amount of time to find a near-optimal solution, and the found solutions are not applicable to new problems in general. To address these challenges, we propose an attention-based reinforcement learning method for the class of job shop scheduling problems by integrating a policy gradient reinforcement learning with a modified transformer architecture. An important result is that our trained learners in the proposed method can be reused to solve large-scale problems not used in training and demonstrate that our approach outperforms the results in recent studies and widely adopted heuristic rules.
Index terms— Combinatorial optimization, Job shop scheduling problem, Reinforcement learning, Attention
1 Introduction
The job shop scheduling problem (JSSP) is a well-known and one of the hardest classes of combinatorial optimization problems in operations research, computer science, and industrial engineering. It has a long history as a research topic and is still being actively studied because of its applicability in numerous industrial areas, such as manufacturing and production domains. (Gao et al.,, 2020).
JSSP has been tackled in mainly two ways, by exact solution approaches and approximate approaches (Zhang et al.,, 2019; Kwon et al.,, 2021; Xie et al.,, 2019). Exact solution approaches with mathematical programming approaches have been widely adopted (Ku and Beck,, 2016; Pei et al.,, 2020; Wang et al.,, 2018; Zhang and Wang,, 2018), however, it may take a prohibitive amount of time to find an optimal solution, or the problems can be infeasible to find an optimal solution. This can be problematic for real-world JSSP, where a large number of jobs or machines would be associated. This has led researchers and practitioners to turn their attention to approximate methods, which aim to achieve near-optimal solutions instead. Various heuristic approaches have been applied to solve JSSP.(Gonçalves et al.,, 2005; Park et al.,, 2003; Peng et al.,, 2015; Kawaguchi and Fukuyama,, 2016; Chakraborty and Bhowmik,, 2015; Cruz-Chávez et al.,, 2017; Nagata and Ono,, 2018; Shi et al.,, 2020)
Although the above mentioned approaches have gained success in addressing JSSP and are recognized as effective means to solve this particular problem type, solutions by those methods are subject to the constraints that are assumed when solving the problems. For example, solutions by the above methods are for the particular problems where the methods are applied, and thus, the solutions are not applicable to other problems in general. New problems are solved anew even with the same approach. Considering the classical solution approaches’ shortcomings it is essential to have a framework that can effectively solve large-scale problems without the need for model adjustments or human intervention.
To address those challenges, we propose a learning method that can be reused to provide solutions to unseen problems. In particular, we employ attention mechanisms (Bahdanau et al.,, 2014) and integrate them with reinforcement learning (Sutton and Barto,, 2018) and the transformer architecture (Vaswani et al.,, 2017). We show that with our method, trained learners can be reused to solve large-scale problems not used in training and demonstrate that our proposed approach outperforms the results in recent studies and widely adopted heuristic rules.
The remainder of the paper proceeds as follows. Related work is provided in Section 2, the problem is defined in Section 3, details of our methodologies are proposed in Section 4, the experiments are described, and the results are provided in Section 5, and the study is concluded in Section 6.
2 Related Work
Vinyals et al. (Vinyals et al.,, 2015) proposed the Pointer Network, a sequence-to-sequence model to generate a permutation of the input sequence instead a generated sequence from a fixed-size dictionary. This allows the model to select and order elements from the input sequence. In the original work of Vinyals et al. (Vinyals et al.,, 2015), the traveling salesman problem (TSP) was studied. However, the model is trained with supervised learning and optimal solutions obtained from alternative methods, a notable limitation is which may not be practical in all cases. To address this issue Bello et al. (Bello et al.,, 2017) introduced a Reinforcement Learning (RL) framework as an alternative to supervised learning to train models. Similar to Vinyals et al. (Vinyals et al.,, 2015), they also employed the Pointer Network architecture and studied the TSP in their research. Kool et al. (Kool et al.,, 2018) applied a transformer architecture Vaswani et al., (2017), with a transformer encoder and a point network type of decoder to solve TSPs and vehicle routing problems (VRPs). Nazari et al. (Nazari et al.,, 2018) extended the VRP, modified the encoder by replacing the LSTM unit with 1D convolutional layers to have the model invariant to the order of input sequence, handling the dynamic state changes.
Subsequent studies followed the RL framework with different architectures and training methods to solve CO problems, mainly for TSPs and VRPS (Miki et al.,, 2018; Deudon et al.,, 2018; Ma et al.,, 2019; Kwon et al.,, 2020; Xu et al.,, 2021). Beginning with well-known TSP and VRP problms, research efforts have extended to the JSSP with a similar framework.
Zhang et al. Zhang et al., (2020) proposed a deep reinforcement learning (DRL) model to learn how to prioritize an operation within a possible operation set. They utilized a graph neural network to embed nodes (operations) in a disjunctive graph Adams et al., (1988) to represent states effectively. They designed the reward function as the quality difference between the partial solutions corresponding to the two states $s_{t}$ and $s_{t+1}$. Rather than employing the common makespan of a schedule as the quality metric, this approach utilizes an estimated completion time to measure the quality of a specific state $s_{t}$. Furthermore, the quality difference between the two states tends to promote a greedy selection by the temporal reward. However, opting for an action with a slightly lower immediate reward can often lead to a more favorable overall schedule. This limitation results in a relatively large optimality gap, $17.7\%$ in even small 6 machines $\times$ 6 jobs synthetic instances. To address these limitations, we use actual makespan as the reward metric and episodic reward instead of temporal reward to prevent greedy selection, We demonstrate better performance, $4.8\%$ optimality gap in $6\times 6$.
Numerous studies Park et al., 2021b ; Park et al., 2021a ; Chen et al., (2022); Yuan et al., (2023) have adopted a graph neural network or its variant to embed operations similar to the one proposed by Zhang et al. (Zhang et al.,, 2020) with different reward functions, training algorithms, masking technique and model structures. The graph neural network shares information between operations connected by conjunctive and disjunctive edges within a disjunctive graph, blending two types of information.
Apart from using a graph neural network, Tassel et al. (Tassel et al.,, 2021) proposed an RL framework with seven designed scalers to capture the current state of individual jobs accurately. The major drawback of the proposed framework is problem size-dependent structure. They have to retrain or make adjustments to solve instances with different scales from the initial training. An advantage of our model is an agnostic problem size from an embedding scheme based on a modified transformer(Vaswani et al.,, 2017) structure.
Chien-Liang et al. Liu et al., (2020) introduced another RL framework based on a convolutional neural network (CNN) architecture. They utilize three matrices to represent states effectively. This study’s limitation lies in its relatively simple action space, primarily comprising several dispatching rules such as FIFO and SPT, which restrict candidates of the operations that can be scheduled and diminish solution quality.
3 Problem Definition
A Job Shop Scheduling Problem (JSSP) contains a set of $J$ jobs, for $j=1,2,\ldots,J$ which are to be processed in a shop that consists of $M$ machines for $m=1,2,\ldots,M$. Each job is comprised of $M$ operations, $O_{jm}$ for for $j=1,2,\ldots,J$ and $m=1,2,\ldots,M$ where each operation $O_{jm}$ can be processed on specific machine $m$. Additionally, each operation $O_{jm}$ has a predefined order $t$ in the sequence of operations. That is, $O^{t_{1}}_{jm}$ can be processed after all the prior operations, $O^{t}_{jm}$ where $t<t_{1}$, has been completed. One machine processes exactly one operation at a time without interruption. The time to complete a operation $O_{jm}$ is called the processing time, denoted as $p_{jm}$. Completion of all the element operations completes a job. A valid sequence of job assignments to machines generates a feasible schedule. The objective is to learn a schedule or a scheduling policy that optimizes a function of feasible schedules, e.g., the minimization of makespan or the maximization of machine utilization.
4 Attention-based Reinforcement Learning for JSSP
Our attention-based, reinforcement learning scheduler (ARLS) has an encoder-decoder structure that leverages feature embedding in a higher dimensional space. The input vector of each operation is encoded by linear projection, which is then transformed into embedding by the encoder using a self-attention mechanism. However, the encoder structure of ARLS is different from that of the original transformer (Vaswani et al.,, 2017).
4.1 Input
The input vectors $\mathbf{o}_{n}$, for $n=1,2,...,J\times M+1$, are concatenated from four vectors: 1) job index, 2) operation order in the job sequence, 3) machine index of the operation, and 4) processing time of the operation. The operation index $n=J\times M+1$ accounts for the cases when no operation is assigned (machine is idled).
The operations are initially encoded via the linear projection of the input as shown in 1.
$\bm{h}^{(0)}=\mathbf{W}^{(0)}\bm{o}_{n}+\bm{b}^{(0)}$ (1)
where $\bm{o}_{n}$, for $n=1,2,...,JM+1$, are the input vectors, $\mathbf{W}^{(0)}$ is the weight matrix, and $\bm{b}^{(0)}$ is the bias vector.
4.2 Encoder
The encoder takes the encoding $\bm{h}^{(0)}$ as input and transforms it to the embedding $\bm{h}^{(L)}$ through multiple, e.g., $L$, encoder layers. We generally follow the structure of the encoder layers suggested by  (Vaswani et al.,, 2017) and introduce novel modifications for the JSSP. A masked, multi-head attention mechanism is utilized in the encoder, unlike the original transformer where simple multi-head attention is used in the encoder. The objective is to have attention better focused on associated operations, e.g., the operations in the same job sequence or the operations for the same machine.
Consider the scaled dot-product attention function
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\intercal}}{\sqrt{d_{K}}}\right)V$ (2)
where $Q,K,V$ are each head’s query, key, and value matrices, respectively, and $d_{K}$ is the dimension for each head’s query and key matrices. Let each element $c_{i,j}=\bm{q}_{i}\bm{k}^{T}_{j}/\sqrt{d_{K}}$ where $\bm{q}$, $\bm{k}$ are $i$th and $j$th columns of the query and key matrices, $Q$ and $K$, respectively.
We apply two different masking schemes to $c_{i,j}$: one related to the jobs, denoted as $c_{i,j}^{\mathcal{(J)}}$, and another for the machines, $c_{i,j}^{\mathcal{(M)}}$, and integrate the two encoder results at the top of the stack. In particular, the values of the elements $c_{i,j}^{\mathcal{(J)}}=-\infty$ are masked if operations $i$ and $j$ are not in the same job sequence, and, similarly, $c_{i,j}^{\mathcal{(M)}}=-\infty$ if operations $i$ and $j$ are not processed by the same machine.
The embedding of operations from the last layer of the encoder $\bm{h}^{(L)}$ is computed as a convex combination of the two embeddings $\bm{h}^{(L)}_{\mathcal{J}}$ and $\bm{h}^{(L)}_{\mathcal{M}}$ as
$\bm{h}^{(L)}=\lambda\bm{h}^{(L)}_{\mathcal{J}}+(1-\lambda)\bm{h}^{(L)}_{\mathcal{M}}$ (3)
where $0\leq\lambda\leq 1$ controls the importance of each embedding.
4.3 Decoder
The decoder has a single pillar of $L^{*}$ identical attention layers similar to the original transformer model  (Vaswani et al.,, 2017) However, masking is performed respective to the operations that were already assigned. Let $\mathcal{O}_{t}^{-}\subset\mathcal{O}$ be the set of operations that were already assigned by time $t$. The complement is $\mathcal{O}_{t}^{+}=\mathcal{O}\setminus\mathcal{O}_{t}^{-}$. Also, $\bm{h}_{t-1}$ denotes the encoder embedding of the operation assigned at the previous time step $t-1$.
For decoder attention, which is also a masked, multi-head attention with scaled dot-product, we use the encoder embedding of the last assigned operation $\bm{h}_{t-1}$ for query vector, $\bm{q}$, and the encoder embeddings from the previous time step $\bm{h}^{(L)}$ for keys $K$ and values $V$
An action selection module generates a probability distribution over possible operations on the top of the decoder. Attention scores are calculated using the decoder output at time $t$, denoted as $\bm{q}_{(t)}$ for query, and the previous encoder embeddings $\bm{h}^{(L)}$ for keys $K$ to generate probabilities $p_{o}$ over the possible operations $\mathcal{O}^{+}$. Nonlinear activation function $tanh$ is applied with the clipping technique (Li,, 2023) before the softmax function is applied.
$p_{o}$ =$=\begin{cases}\text{softmax}\left(C\text{tanh}\left(\bm{q}_{(t)}\bm{k}_{o}^{\intercal}/\sqrt{d_{K}}\right)\right),&\text{if }\text{operation}\,\,j\notin\mathcal{O}^{-}\\ 0&\text{otherwise}\end{cases}$ (4)
where $C$ is the clipping constant. The probability $p_{o}$ is used to choose operation at time step $t$.
4.4 Training
We modify the Monte-Carlo policy gradient method (Sutton and Barto,, 2018; Bahdanau et al.,, 2016) for training. The performance of learned policy was estimated by a Monte-Carlo method using observed makespans. The average of the makespans respective to the same problem is used as the baseline. Observed makespans minus the baseline, generate the advantages used for gradient updates
$\nabla_{\theta}J(\pi_{\theta})=\underset{\tau\sim\pi_{\theta}}{\mathrm{E}}\left[\sum_{t=0}^{T}\nabla_{\theta}\ln\pi_{\theta}(a_{t}|s_{t})\left(R(\tau)-b\right)\right]$ (5)
where $\tau$ is a trajectory of an episode, $R(\tau)$ is the Monte-Carlo episodic reward (makespan of a trajectory), and $b$ is a baseline to reduce variance. The value from the value network is used as $b$ in many studies Mnih et al., (2013); Schulman et al., (2015); Haarnoja et al., (2018); Reda et al., (2020).
A different strategy compared to typical Monte-Carlo Policy Gradient generates multiple trajectories. A multi-trajectory strategy offers several advantages compared to employing a value network. More problem instances are expected to increase the scheduling performance robustness. And for each problem instance, the learner is expected to improve with more diverse scheduling solutions.
Let $\mathcal{T}$ denote the set of all trajectories, that is, all the feasible solutions for a JSSP instance. Let each $\tau_{n}\in\mathcal{T}$ be a sampled trajectory from time $t=0$ to $T$ where $T$ is the time a trajectory completes a schedule. For simplicity, we ignore the fact that $T$ can vary. and $\mathcal{T}=\langle\tau_{1},\ldots,\tau_{N}\rangle$ be the set of $N$ sampled trajectories for a JSSP instance. We sample $N$ parallel trajectories that enable us to produce a batch of disparate trajectories in parallel and enrich feasible solutions obtained for individual problem instances. One can control the sample size $N$. The equation 5 will be replaced with the following equation 6 for parallel trajectory.
$\nabla_{\theta}J(\pi_{\theta})=\frac{1}{N}\sum_{n=1}^{N}\left(\underset{\tau\sim\pi_{\theta}}{\mathrm{E}}\left[\sum_{t=0}^{T}\nabla_{\theta}\ln\pi_{\theta}(a_{t}|s_{t})\left(R(\tau_{n})-b\right)\right]\right)$ (6)
where $R(\tau_{n})$ is the Monte-Carlo episodic reward of trajectory $n$, and $b=\frac{1}{N}\sum_{n=1}^{N}R(\tau_{n})$, which is the average of the $N$ episodic rewards obtained from the same problem instance in parallel.
5 Experimental Results
The training and testing approach for ARLS prioritizes rapid model learning in simplified settings. A key objective is a learner design which effectively scales to greater problem instance sizes and maintains high performance. ARLS is trained with synthetic data with 6 jobs and 6 machines. In total, 36 operations, with processing times randomly selected uniformly in the range between 1 and 15, similar to the prior study Zhang et al., (2020). The Adam optimization (Kingma and Ba,, 2014) was employed for weight updates. ARLS uses a relatively low complexity transformer architecture with only three encoder layers and a single decoder layer, the model dimension is 256, the number of heads is 16, and the fully connected feed-forward neural networks dimension is 512. Further design studies might improve performance.
Training generated 100,000 replicate synthetic 6 jobs $\times$ 6 machines problem instances. A mini-batch  (Yang et al.,, 2019) update is applied. The objective is to minimize the schedule’s makespan. To evaluate ARLS performance, the average optimality gap is used
[…]
ARLS generalized performance is tested at varying sizes of synthetic and benchmark instances, including problem instance sizes greater than seen in training. To evaluate, we use three synthetic datasets with sizes $6\times 6$, $10\times 10$ and $15\times 15$. Additionally, seven well-known benchmark sets, from Shylo, (2010), are tested for performance.
Table 1 shows the test results on synthetic data. The average optimality gap is obtained from 1,000 instances for each problem size, which is compared with previous studies by Z20 (Zhang et al.,, 2020). Our model is trained on only $6\times 6$ setting. In contrast, Zhang et al., (Zhang et al.,, 2020) trained their model separately with samples from the three problem sizes. On average, ARLS outperforms their results by a substantial margin of 12.5%.
Table 1: AReL Performance on synthetic dataset
Jobs Machines Instances ARLS (Our) Z20
6 6 1000 4.8 17.7
10 10 1000 10.9 22.3
15 15 1000 13.5 26.7
ARLS was tested with seven benchmark datasets of varying sizes based on the number of jobs (ranging from 6 to 100) and machines (ranging from 5 to 20). Table 2 shows ARLS’s performance compared to work by Zhang et al., (2020) denoted as Z20, Park et al., 2021a as P21a, Park et al., 2021b as P21b, Chen et al., (2022) as C22, and Yuan et al., (2023) and Y23, and popularly adopted heuristic rules, such as first-in-first-out (FIFO), shortest processing time (SPT), and most work remaining (MWKR). Here, instead of the optimality gap, the best known gap, which is the relative difference to the best known makespan is used (as optimal makespans are unknown for some of the problems). Problems with multiple instances show average results. Our ARLS model showed the best performance on all but 4 problem sizes of 26 problem classes, despite being trained in the simplest setting of 6x6.
Table 2: Performance of our ARLS model compared with heuristic rules and other studies on benchmark datasets. Performance is evaluated by the best known gap in percentage (with averages for multiple instances). Results by Zhang et al., (2020) denoted as Z20, Park et al., 2021a as P21a, Park et al., 2021b as P21b, Chen et al., (2022) as C22, and Yuan et al., (2023) and Y23, and popularly adopted heuristic rules, such as first-in-first-out (FIFO), shortest processing time (SPT), and most work remaining (MWKR).
Dataset Jobs Machines Instances FIFO SPT MWKR Z20 P21a P21b C22 Y23 Ours
ABZ 10 10 2 14.8 12.9 9.5 - 6.2 10.1 - 7.1 6.1
15 15 3 31.3 33.3 21.3 - 20.5 29.0 - 22.3 17.6
FT 6 6 1 18.2 60.0 10.9 - 7.3 29.1 - 9.1 7.3
15 15 1 27.3 15.5 24.9 - 19.5 22.8 - 18.5 11.3
20 5 1 41.2 8.8 34.8 - 28.7 14.8 - 14.3 10.1
YN 20 20 4 25.9 30.6 20.8 - 18.4 24.8 - 20.3 16.1
SWV 20 10 5 44.4 26.3 38.7 - 34.4 28.4 - 24.8 21.7
20 15 5 44.9 32.0 36.7 - 30.5 29.4 - 28.3 21.3
50 10 10 30.0 17.0 25.6 - 25.3 16.8 - 14.1 10.3
ORB 10 10 10 29.7 26.3 29.9 - 20.0 21.8 - 20.9 13.5
TA 15 15 10 21.2 25.9 23.1 26.0 15.3 20.1 35.4 17.9 13.1
20 15 10 22.8 32.8 30.0 30.0 19.5 25.0 32.1 20.6 17.9
20 20 10 24.0 27.8 27.7 31.6 17.2 29.3 28.3 19.5 16.1
30 15 10 24.0 35.3 30.2 33.0 19.1 24.7 36.4 22.4 19.6
30 20 10 25.7 34.4 30.9 33.6 23.7 32.0 34.7 23.3 22.8
50 15 10 17.8 24.1 20.1 22.4 13.9 15.9 31.9 16.0 12.9
50 20 10 18.4 25.2 23.2 26.5 13.5 21.3 28.0 17.4 16.8
100 20 10 8.8 14.4 12.8 13.6 6.7 9.2 18.0 8.9 9.3
DMU 20 15 10 30.5 64.1 37.2 39.0 - - - 26.6 18.9
20 20 10 26.4 64.6 32.4 37.7 - - - 23.6 17.3
30 15 10 34.8 62.6 39.3 41.9 - - - 28.7 20.7
30 20 10 32.2 65.9 36.6 39.5 - - - 29.2 22.6
40 15 10 31.2 55.9 35.1 35.4 - - - 24.5 20.2
40 20 10 33.2 63.0 39.7 39.4 - - - 30.1 24.8
50 15 10 31.0 50.4 34.7 36.2 - - - 24.9 18.8
50 20 10 35.3 62.2 41.4 38.4 - - - 29.5 21.8
Figure 1: Our ARLS model’s performance on benchmark datasets, ABZ, FT, YN, SWVM, and ORB
On average, ARLS outperformed other algorithms all cases, as shown in figures 1 and 2. Only in a few problem classes in TA was ARLS exceeded. Even then, the P21a algorithm was trained on problem instances more that 100% greater than the training size for ARLS.
Figure 2: Our ARLS model’s performance on benchmark datasets TA and DMU
6 Conclusions
We developed an attention-based reinforcement learner for job shop scheduling problems (ARLS) with a modified transformer architecture. A key point was o enable the learner to scale with problem instance size. The learner was trained with synthetic problem instances of a relatively small size in the number of jobs and machines and scaled to new problems of various sizes with both synthetic and benchmark datasets. We have obtained results in general outperforming results in previous studies in the experiments and compared to widely adopted heuristic rules. The model architecture is new, but the transformer is low in complexity (few layers), and further work might improve results with larger transformer architectures.